Is the formula an = -4n(n-1) explicit or recursive? Find the first five terms of the sequence.
Answer by jim_thompson5910(35256)

It's explicit because you can find ANY term without having to know anything about the previous term (or terms).

First term:

an = -4n(n-1)

a1 = -4*1*(1-1)

a1 = -4*0

a1 = 0
-------------------------------------------------------

Second term:

an = -4n(n-1)

a2 = -4*2*(2-1)

a1 = -8*1

a1 = -8

I'll let you finish.